### v1

# SCINE — Software for Chemical Interaction Networks

Abstract
The software for chemical interaction networks (SCINE) project aims at pushing the frontier of quantum chemical calculations on molecular structures to a new level. While calculations on individual structures as well as on simple relations between them (e.g., as given by an intrinsic reaction coordinate) have become routine in chemistry, new developments have pushed the frontier in the field to high-throughput calculations. Chemical relations may be created by a search for specific molecular properties in a molecular design attempt or they can be defined by a set of elementary reaction steps that form a chemical reaction network. The software modules of SCINE have been designed to facilitate such studies. The features of the modules are (i) general applicability of the applied methodologies ranging from electronic structure (no restriction to specific elements of the periodic table) to microkinetic modeling (with little restrictions on molecularity), full modularity so that SCINE modules can also be applied as stand-alone programs or be exchanged for external software packages that fulfill a similar purpose (to increase options for computational campaigns and to provide alternatives in case of tasks that are hard or impossible to accomplish with certain programs), (ii) high stability and autonomous operations so that control and steering by an operator is as easy as possible, and (iii) easy embedding into complex heterogeneous environments for molecular structures taken individually or in the context of a reaction network. A graphical user interface unites all modules and ensures interoperability. All components of the software have been made available open source and free of charge.
preprint: AIP/123-QED
I Introduction
The past decades have seen a steady rise in computing power. Vast computational resources have become available, allowing for large-scale computing campaigns such as virtual high-throughput screeningShoichet (2004); Pyzer-Knapp et al. (2015). Hence, more challenging tasks have become possible of which the exploration of complex chemical reaction networks (CRNs) is one exampleDewyer, Argüelles, and Zimmerman (2018); Vázquez, Otero, and Martinez-Nunez (2018); Simm, Vaucher, and Reiher (2019); Maeda and Harabuchi (2021); Baiardi et al. (2022). However, the complexity of these tasks implies that their automation is not straightforward. In fact, while standard quantum chemical software provides building blocks, these need to be well integrated into a meta-algorithm that orchestrates and steers the computational campaign to allow for the automated execution of complex exploration procedures with only little human input. Hence, a new kind of software for such a purpose is needed, which our software for chemical interaction networks (SCINE) project aims to provide. Under the umbrella of SCINE, a large number of different software packages is being developed. The primary focus of SCINE, uniting all these modules, is to provide software for predictive quantum chemistry.
The first-principles prediction of reaction mechanisms in chemistry encompasses the ability to find all possible reactions between a set of reactants, no matter what these reactants are. Therefore, one has to be able to cope with complicated electronic structures (such as molecules with multiconfigurational character), very large molecules or molecular aggregates (such as molecules adsorbed on a surface), and solvation effects, to name only a few challenges. As a consequence, a software stack allowing to study a chemical system from a truly holistic point of view needs to provide broad and stable functionality. This, in turn, results in challenges concerning the flexibility, ease of use, and long-term maintainability of the software.
Within SCINE, we address these challenges by a strictly modular architecture, a loose coupling between the individual modules, and consistent interfaces between them. Individual modules which are only loosely coupled in terms of their source code (i.e., the flow of data between them is clearly defined and limited to only a few instances) are easier to maintain than a complex monolithic architecture. At the same time, this makes developing the software easier and less error-prone, as so-called side effects (i.e., unexpected and unwanted effects upon changing the source code) are largely minimized. Finally, this approach naturally leads to individual modules that are fully functional programs in their own right with useful features. For less complex tasks, it is therefore sufficient to install only a single or a few required modules without having to deal with a particularly complicated software stack.
Complex computational procedures will require many SCINE modules to work together. Well-defined interfaces between these modules largely simplify the creation of workflows, leveraging the functionality provided by different modules. We strive to make these interfaces as well as the overall architecture as similar and consistent as possible across the entire SCINE software stack. Therefore, different modules provide a similar appearance in such a way that someone experienced with the code of a particular module should quickly be able to familiarize themselves with any other module. This standardization of the code is, therefore, of crucial importance to reduce the complexity of working with it.
Finally, we believe that the entire functionality provided by SCINE should be freely available to the academic community. Therefore, we provide all the modules under the permissive three-clause BSD license.
Figure 1: An overview of the SCINE modules arranged according to task classes.
Fig. 1 presents a high-level overview of the SCINE modules presented in this work. In this figure, the modules are grouped according to the tasks they are responsible for. Many modules (for example, Sparrow, Swoose, and Parrot; also many external programs such as ORCA and Turbomole are supported) deliver information on potential energy surfaces (PES), i.e., they calculate electronic energies, nuclear gradients, Hessian matrices, and so on. These quantities are essential inputs for ReaDuct, which is devoted to structure manipulation tasks such as transition state searches. Even though it can be operated in a stand-alone manner, in the context of a high-throughput screening or exploration of a chemical reaction network, a large number of calculations needs to be carried out. For this, we have the calculation backend Puffin. Puffin relies on a database to store all calculation inputs and outputs. Chemoton makes use of these data to drive the exploration of CRNs. Once a network has been established, the module KiNetX can establish a microkinetic model to analyze the concentration fluxes through this network (such a microkinetic analysis can even be applied during a running calculation; see below). Many SCINE modules operate on a graph-based representation of molecular structures; for this we developed the module Molassembler. Finally, there is a graphical user interface (GUI), called Heron, which eases interacting with SCINE modules.
This work is structured as follows: In section II, we will first give an in-depth overview of important SCINE modules and the data flows between them. Afterwards, every module will be presented in more detail in its own subsection. In these subsections, we will also highlight similar software developments by others whenever appropriate.
II Overview of SCINE Modules
A high-level overview of most SCINE modules and of the data flows connecting them is shown in Fig. 2.
Figure 2: SCINE modules and their communication.
ChemotonBensberg et al. (2023a); Unsleber, Grimmel, and Reiher (2022); Simm and Reiher (2017) is the central module for the exploration of chemical reaction networks. During an exploration, it keeps track of all compounds and reactions found so far (for this bookkeeping, it relies on functionality for the analysis of chemical structures provided by MolassemblerBensberg et al. (2023b); Sobez and Reiher (2020), see below), properly stores all these data in a database, and sets up new calculations to drive the exploration forward. Chemoton does not carry out these calculations but writes all necessary input data to the database.
The calculations are then carried out by another module called PuffinBensberg et al. (2023c). Puffin retrieves information on new calculations from the database, carries them out, and then writes the results back to the database. For carrying out a calculation, Puffin typically invokes other SCINE modules. For example, it can call SwooseBrunken, Csizi, and Reiher (2021); Brunken and Reiher (2020, 2021) for QM/MM calculations or KiNetXBensberg et al. (2023d); Proppe and Reiher (2019); Bensberg and Reiher (2023a) for microkinetic modeling. For most calculations, however, Puffin relies on ReaDuctBensberg et al. (2023e). The ReaDuct module carries out a wide range of structure-manipulating tasks such as structure minimizations, transition state searches, intrinsic reaction coordinate calculations, B-spline interpolation and optimization of reaction pathsVaucher and Reiher (2018), and Newton trajectory calculationsUnsleber, Grimmel, and Reiher (2022).
For these calculations, ReaDuct in turn relies on quantum chemistry software, which delivers properties such as electronic energies, nuclear gradients, and Hessian matrices. ReaDuct interfaces to many external program packages. AMSte Velde et al. (2001), CP2KKühne et al. (2020), DFTB+Hourahine et al. (2020), GaussianFrisch et al. , MRCCKállay et al. (2020), ORCANeese (2022), TurbomoleFranzke et al. (2023), SerenityUnsleber et al. (2018); Niemeyer et al. (2022), and xTBBannwarth et al. (2021) are currently supported.
Apart from these external software packages, ReaDuct is also interfaced to SparrowBosia et al. (2023a); Husch, Vaucher, and Reiher (2018); Bosia, Weymuth, and Reiher (2022), which is a SCINE module providing many semi-empirical methods, such as AM1, PM6, DFTB3, and also the OMx methodsBosia et al. (2023b). Moreover, Sparrow is the backend for interactive quantum chemical calculations (see below).
Molassembler is a general-purpose cheminformatics toolkit for the modeling of organic and inorganic moleculesSobez and Reiher (2020). It represents a molecule as an annotated graph which includes information about the local coordination geometries of the atoms and stereochemical information. During an exploration, it is a key ingredient to compare structures. Since Molassembler can also create three-dimensional coordinates from its graph representation, it can be used to create conformers of a given molecule.
Large-scale explorations typically require millions of quantum chemical calculations. Therefore, fast methods with sufficient accuracy are of vital importance. Machine learning potentials can offer both high accuracy and high speed when trained on sufficiently accurate data. However, such potentials require extensive and costly training and often have difficulties learning additional data without “forgetting” previous knowledge. Moreover, most structural descriptors cannot efficiently represent many different atom types. We recently addressed these problems by introducing the new concept of lifelong machine learning potentialsEckhoff and Reiher (2023) implemented in the SCINE module Parrot. Other well-known machine learning potentials such as ANISmith, Isayev, and Roitberg (2017); Devereux et al. (2020) and M3GNetChen and Ong (2022) are also available in Parrot.
Since Chemoton is based on the first principles of quantum mechanics, it is applicable to any chemical system. In practice, this might require specific quantum chemical methods. For example, for very large systems such as proteins, even fast semiempirical methods like the ones provided by Sparrow are typically too time-intensive for large-scale explorations. In this case, one can resort to molecular mechanics calculations. To enable such type of modeling out of the box, the SCINE software stack includes a module called Swoose. Swoose can provide energies and forces from molecular mechanics for any chemical system and run molecular mechanics simulations. If no parameters are available for the system, it can automatically parametrize a system-focused force field from quantum chemical reference calculationsBrunken and Reiher (2020). Large systems for which accurate reference calculations would be too costly are automatically fragmented. Moreover, Swoose is also able to carry out QM/MM calculations, in which a key region within a system is automatically defined and then described by an electronic structure model, while the rest of the system is described by a standard force field (supported by machine learning corrections)Brunken and Reiher (2021).
Open-shell systems often require a multi-configurational wavefunction ansatz to be described qualitatively correctly. Multi-configurational methods such as the density renormalization group (DMRG)White (1992, 1993); Chan et al. (2008); Chan and Zgid (2009); Marti and Reiher (2010); Schollwöck (2011); Chan and Sharma (2011); Wouters and Van Neck (2013); Kurashige (2014); Olivares-Amaya et al. (2015); Szalay et al. (2015); Yanai et al. (2015); Baiardi and Reiher (2020) typically require the definition of a so-called active orbital space on input. However, this active space is often not trivial to determine. Our autoCAS algorithmStein and Reiher (2016); Stein, von Burg, and Reiher (2016); Stein and Reiher (2017), implemented in a module of the same nameStein and Reiher (2019); Bensberg et al. (2023f), provides an automated way to reliably determine the active space of any molecule. Moreover, once the active space is determined, autoCAS can also invoke software to carry out multi-configurational calculations. Besides QCMaquisKeller et al. (2015), which is our software for DMRG calculations and driven by autoCAS via OpenMolcasLi Manni et al. (2023), autoCAS is also interfaced to ChronusQWilliams-Young et al. (2020), to PySCFSun et al. (2020), and SerenityUnsleber et al. (2018). We note that QCMaquis can do much more than just calculate the ground state electronic energy of a molecule. It can also optimize excited statesBaiardi et al. (2019); Baiardi, Kelemen, and Reiher (2022), solve the vibrational Schrödinger equation for anharmonic potential energy surfacesBaiardi et al. (2017); Glaser, Baiardi, and Reiher (2023), and carry out quantum dynamics simulationsBaiardi and Reiher (2019); Baiardi (2021). Furthermore, it features embedding approachesDresselhaus et al. (2015); Hedegård and Reiher (2016), a non-orthogonal state interaction methodKnecht et al. (2016), relativistic calculationsBattaglia, Keller, and Knecht (2018), multi-component DMRGMuolo et al. (2020); Feldmann et al. (2022), and N-electron valence state perturbation theory calculaitonsFreitag et al. (2017).
A central module is HeronMüller et al. (2024); Bensberg et al. (2022), our graphical user interface. It aims to make the entire SCINE software ecosystem easily accessible. Heron provides the ability to browse the contents of a database, visualize and analyze existing reaction networks, and operate Chemoton (for example, with the so-called Steering WheelSteiner and Reiher (2023)). Moreover, it is a graphical user interface for Swoose, ReaDuct, autoCAS, and Sparrow. Together with the latter, it provides the ability to interactively carry out quantum chemical calculations, in which the response of a molecule to structural manipulations can be experienced in real time. This can even be coupled with haptic devices, allowing a person to literally feel the forces acting on atoms in a moleculeMarti and Reiher (2009); Haag and Reiher (2013); Haag et al. (2014); Weymuth and Reiher (2021).
For the sake of completeness, we mention two modules that are essential for the entire SCINE software stack, but they are typically not very visible when carrying out calculations. The first of these modules is CoreBosia et al. (2023c), which provides the common so-called Calculator interface. This interface is crucially important for the high interconnectivity yet loose coupling of the individual SCINE modules. The second module is called UtilitiesBaiardi et al. (2023); it contains much functionality needed by many of the other modules. For example, it contains optimization algorithms, data structures for objects such as atomic orbitals and density matrices, functions to read and write molecular coordinates to standard file formats such as XYZ and PDB, and so forth.
In the following, we present SCINE modules in more detail. For each module, we describe the field of application, its input and output, as well as its application programming interface (API), and we illustrate the latter with code snippets.
II.1 Molassembler
II.1.1 Field of application
MolassemblerBensberg et al. (2023b); Sobez and Reiher (2020) provides characterization, manipulation, and conformer generation for organic and inorganic molecules. In principle, it can handle any organic or inorganic molecule, including haptic and multidentate ligands and almost arbitrary coordination spheres. Its general approach to the characterization of molecules and its broad applicability makes it distinct compared to similar programs such as RDKitRDK or openBabelO’Boyle et al. (2011), which focuses on organic molecules, or molSimplifyIoannidis, Gani, and Kulik (2016); Edholm et al. (2023), AARONGuan et al. (2018) and DENOPTIMFoscato, Venkatraman, and Jensen (2019), which provide utilities to design new molecules, or CRESTPracht, Bohle, and Grimme (2020) for conformer generation.
Molassembler characterizes a molecule’s connectivity by a graph and the local geometry of each atom by polyhedral shapes. In organic chemistry, atoms typically have only up to four bonding partners, leading at most to a tetrahedral local environment. By contrast, transition metal atoms can have significantly more neighbors and diverse coordination geometries. Therefore, Molassembler supports polyhedra for the local geometry up to icosahedra and cuboctahedra and can be applied to the full spectrum of organic and inorganic chemistry. Furthermore, it supports multidentate and haptic ligands. Molassembler differentiates stereoisomers through a ranking algorithm based on generalizing the ranking rules for organic molecules provided by the International Union of Pure and Applied ChemistryFavre and Powell (2013). The abstract molecule representation provided by Molassembler can be manipulated to change the connectivity in the molecule and replace or edit ligands to generate new molecules while retaining any chiral information in the remaining molecule wherever possible. Furthermore, Molassembler can generate conformers by enumerating the rotamers of the molecule, generating initial Cartesian coordinates for the conformer guesses before refining the coordinates.
II.1.2 API
Molassembler can be used through its Python API or as a C++ library. It generates its abstract molecule representation based on either Cartesian coordinates and bonding pattern, its SMILESWeininger (1988) or its InChiHeller et al. (2015) representation. Furthermore, it can read and write its own molecule representation as a JSON document or binary string.
Figure 3: Example iron–phosphorus complex manipulated with Molassembler.
We demonstrate a simple molecule manipulation with Molassembler in the code extract 1. We first create Molassembler’s internal molecule representation of an iron–phosphorus complex by reading its SMILES string in line 5. Then, we create the vector graphic in Fig. 3(a), displaying the molecule, cutting one of the iron-phosphorus bonds in line 8, and adding a water ligand in lines 10 to 12. Finally, we save a vector graphic displaying the final molecule [see Fig. 3(b)].
1import scine_utilities as utils
2from scine_molassembler import io
3
4complex_smiles = "Cl[FeH-4]56([P+](C)(C)CC[P+]5(C)C)[P+](C)(C)CC[P+]6(C)C"
5fe_complex = io.experimental.from_smiles(complex_smiles)
6io.write("mol.svg", fe_complex)
7
8fe_complex.remove_bond(1, 2)
9
10new_atom_index = fe_complex.add_atom(utils.ElementType.O, 1)
11fe_complex.add_atom(utils.ElementType.H, new_atom_index)
12fe_complex.add_atom(utils.ElementType.H, new_atom_index)
13
14io.write("mol2.svg", fe_complex)
Listing 1: Example code for molecule manipulations with Molassembler.
Based on Molassembler’s molecule representation, we can then generate conformers of the molecule, as shown in the code extract 2. In this example, we let Molassembler generate an ensemble of 100 conformers.
1from scine_molassembler import dg
2results = dg.generate_ensemble(fe_complex, 100, seed=42)
Listing 2: Conformer generation with Molassembler.
The conformer generation’s result (results) is a list containing the Cartesian coordinates of the conformers, which could be analyzed further with other tools within the SCINE framework.
II.2 Utilities
II.2.1 Field of application
Certain functionality (e.g., reading in molecular structures from XYZ files) is required in several SCINE modules. Having separate functions for this in every individual module would lead to code which is difficult to maintain. Therefore, we created the separate module SCINE UtilitiesBaiardi et al. (2023) to collect functionality shared across the entire SCINE project. SCINE Utilities is rarely used in a stand-alone manner (even though it could) since the functionality provided by it is typically needed in the more complex tasks and workflows provided by other SCINE modules.
From the wealth of functionality provided by Utilities we shall explicitly mention here data structures to store objects such as molecular structures, atomic orbitals, and density matrices; interfaces to external quantum chemistry programs such as TurbomoleFranzke et al. (2023) and ORCANeese et al. (2020); different optimization algorithms (for example, the well-known L-BFGS algorithmNocedal (1980)); a self-consistent field algorithm including orbital steeringVaucher and Reiher (2017); code for calculating thermodynamic quantities such as the Gibbs free energy relying on the rigid-rotor–harmonic-oscillator model; an implementation of Grimme’s semiclassical D3 dispersion correctionsGrimme et al. (2010); basic molecular dynamics simulation algorithms; functions to handle different chemical file formats (like XYZ and PDB); code for the placement of explicit solvent molecules around a solute; and finally, a lot of technical functionality required across all SCINE modules (for example, data structures and functions to handle calculation settings).
II.2.2 API
The easiest way to interact with the Utilities module is via its Python bindings. As an example, the code in listing 3 first uses data structures provided by Utilities to manually construct a helium dimer with an interatomic distance of 5.0 bohr. Then, this structure is optimized (with the BFGS algorithm as implemented in the Utilities module) by minimizing its energy; the energy is obtained from a simple Lennard-Jones potential, which is also provided by the Utilities module. Finally, the optimized structure is written to an XYZ file.
1import scine_utilities as utils
2
3elements = [utils.ElementType.He, utils.ElementType.He]
4positions = [[0.0, 0.0, 0.0], [5.0, 0.0, 0.0]]
5structure = utils.AtomCollection(elements, positions)
6
7calculator = utils.core.get_calculator("lennardjones")
8calculator.structure = structure
9calculator.set_required_properties([utils.Property.Gradients, utils.Property.Energy])
10
11opt_log = utils.core.Log()
12minimized_structure = utils.geometry_optimize(calculator, opt_log, utils.Optimizer.Bfgs)
13
14utils.io.write("minimized_structure.xyz", minimized_structure)
Listing 3: Simple structure minimization with the Utilities module.
II.3 Calculator Interface
The core of the modular SCINE infrastructure is a custom Calculator interface providing the (abstract) functionality to calculate electronic structure information, such as energy and gradients, for a given set of nuclear coordinates. We have developed interfaces to various programs that allow one to choose methods ranging from molecular mechanics, semi-empirical methods, density functional theory (DFT) to coupled cluster theory. These interfaces will be discussed in more detail in the following sections.
II.4 Sparrow
II.4.1 Field of application
SparrowBosia et al. (2023a); Bosia, Weymuth, and Reiher (2022); Bosia et al. (2023b) specializes in ultra-fast quantum chemical calculations with semi-empirical methods, which are essential in applications such as interactive quantum chemistry or large-scale reaction network exploration. Calculations with semi-empirical methods based on the neglect of differential overlap approximation — NDDO methods — as well as with methods based on the density functional tight binding approach — DFTB methods — are available. Ground state calculations yield electronic energies, Hessian matrices, nuclear gradients, bond orders, and thermodynamic properties such as Gibbs free energies. Electronically excited states can also be calculated — in a configuration interaction approach with NDDO methods or by solving the linear response eigenvalue problem in a time-dependent approach with DFTB methods.
II.4.2 I/O
Sparrow can be called from the command line. In the simplest case, the binary is called with only one command-line argument specifying the path to an XYZ file containing Cartesian coordinates:
1sparrow -x molecule.xyz
Command-line options are available to specify the calculation. For instance, the ground state energy and the Hessian of a doublet water cation with PM6 can be obtained with the command
1sparrow -x h2o.xyz -c 1 -s 2 -M PM6 -H
II.4.3 API
Sparrow is integrated into various SCINE workflows through its Python bindings. Python bindings from SCINE Utilities are required to instantiate a calculator object in the first step. The code snippet in listing 4 shows how the calculator object is created by specifying the desired method. In the second step, the atomic structure is read in from a XYZ file and assigned appropriately. Next, desired (non-default) settings in the form of a dictionary object are defined. Lastly, the desired properties are requested.
1import scine_utilities as utils
2import scine_sparrow as sparrow
3
4calculator = utils.core.get_calculator("pm6",   "sparrow")
5atoms = utils.io.read("h2o.xyz")[0]
6calculator.structure = atoms
7calculator.settings.update({
8    "spin_mode": "unrestricted",
9    "molecular_charge": 1,
10    "spin_multiplicity": 2
11})
12calculator.set_required_properties([utils.Property.Energy, utils.Property.Hessian])
Listing 4: Creating and defining the calculator object.
Once the calculator object is fully specified, the actual calculation is run with the command in Listing 5.
1results = calculator.calculate()
Listing 5: Carrying out the calculation.
The results can be saved into dedicated files or simply printed to the console.
II.5 Swoose
II.5.1 Field of application
If a system of interest is too large for a full quantum mechanical description, a quantum–classical hybrid model must be constructed. Such models generally describe a small subsystem quantum mechanically (QM) and the remainder with a molecular mechanical (MM) model which parametrizes inter- and intramolecular interactions based on experimental or calculated reference data. The module SCINE Swoose provides such MM models, the framework to combine its models with any existing electronic structure theory to construct a QM/MM hybrid model, and workflows specialized for nanoscopic systems. Generally, MM models can be difficult to construct due to the structural complexity of nanoscopic systems and the necessity of MM models to define bonds and charges a priori. Swoose overcomes this hurdle by providing a general workflow for structure preparationCsizi and Reiher (2023) and a system-focused atomistic model (SFAM)Brunken and Reiher (2020) that can be generated for each system of interest with automated QM reference data generation by fragmenting the system, carrying out numerous structure optimizations in parallel, and fitting force constants with a partial Hessian procedure. The generated structural and MM model can then be leveraged easily in automated proceduresCsizi, Steiner, and Reiher (2023). Reaction mechanisms can be studied with hybrid models, for which the QM region can be automatically determined by another algorithm implemented in Swoose.Brunken and Reiher (2021) It generates multiple candidates for the QM region around an atom alongside multiple large reference systems, calculates the first-order partial derivatives for each nucleus, i.e., the atomic forces, and compares them against those in the reference systems. The algorithm assumes that small errors in atomic forces are a good descriptor together with energies as shown in Ref. 16. It then selects the smallest QM system that features errors below a given threshold. Compared to other approaches to determine a QM regionSumner, Söderhjelm, and Ryde (2013); Liao and Thiel (2013); Kulik et al. (2016); Karelina and Kulik (2017); Brandt and Jacob (2023a, b), to determine protonation statesDolinsky et al. (2004, 2007); O’Boyle et al. (2011); Olsson et al. (2011); Anandakrishnan, Aguilar, and Onufriev (2012); Sastry et al. (2013); Riojas and Wilson (2014); Bochevarov et al. (2016); Eastman et al. (2017); Reis et al. (2020), and to fit a system specific modelGrimme (2014); Vanduyfhuys et al. (2015); Welsh and Allison (2019); Behler and Csányi (2021); Friederich et al. (2021); Unke et al. (2021); Deringer et al. (2021); Musil et al. (2021), Swoose exploits the straightforward parallelization of multiple calculations within the SCINE framework and provides a complete pipeline from experimental data to systematic reaction explorations with QM/MM hybrid models, including real-time calculationsCsizi, Steiner, and Reiher (2023).
II.5.2 I/O
The algorithms for structure preparation, model generation, and quantum region selection can be carried out with Swoose with an input file following the YAML syntax, which allows one to specify all required settings in a straightforward way. Additionally, Swoose can carry out single-point energy calculations directly, structure optimizations, and molecular dynamics simulations on the command line.
II.5.3 API
The Python bindings of the Swoose module provide the same functionality as the executable. The Python module consists of functions that take a single structure file and optional keyword arguments. Additionally, it includes the Parametrizer and QmRegionSelector classes (which can carry out the parametrization and quantum region selection tasks), gives access to the task-specific settings and to additional data structures related to these tasks.
II.6 Parrot
II.6.1 Field of application
Machine learning potentials offer a way to combine accuracy and efficiency in energy and gradient calculations. Their training is typically based on accurate data from electronic structure calculations. Machine learning potentials can preserve this accuracy, while their evaluation inflicts little computational demands comparable to those of force fields. In this way, machine learning potentials can enable accurate high-throughput screening. Parrot provides an interface to machine learning potential predictions of energy and gradients for a chemical structure defined by its elements and atomic positions. In addition, Parrot enables numerical Hessian calculations. Atomic charges and bond orders are obtained by GFN2-xTBBannwarth, Ehlert, and Grimme (2019) if the machine learning method is not able to predict these data.
II.6.2 I/O
Parrot interfaces APIs of established pretrained machine learning potential methods (currently TorchANIGao et al. (2020), Materials Graph LibraryKo et al. (2023), and MACEBatatia et al. (2022)). Reading of model parameter files is handled by the APIs of the respective machine learning potential. Moreover, Parrot allows for reading and predictions of lifelong machine learning potentialsEckhoff and Reiher (2023). Training and saving of improved lifelong machine learning potentials is not yet available in Parrot.
II.6.3 API
Parrot enables the selection of machine learning potentials as potential-energy models in SCINE in the same way as it is the case for other methods such as DFT, semi-empirical methods, and so forth. An overview of all machine learning potentials available in Parrot is provided in Table 1.
Table 1: Machine learning potentials available in Parrot.
Method Family Method
ani Gao et al. (2020) ani2x Devereux et al. (2020)
ani Gao et al. (2020) ani1ccx Smith et al. (2019)
ani Gao et al. (2020) ani1x Smith et al. (2018)
m3gnet Ko et al. (2023) m3gnet-mp-2021.2.8-pes Chen and Ong (2022)
m3gnet Ko et al. (2023) m3gnet-mp-2021.2.8-direct-pes Chen and Ong (2022)
mace Batatia et al. (2022) mace-mp_large Batatia et al. (2023)
mace Batatia et al. (2022) mace-mp_medium Batatia et al. (2023)
mace Batatia et al. (2022) mace-mp_small Batatia et al. (2023)
mace Batatia et al. (2022) mace-off_large Kovács et al. (2023)
mace Batatia et al. (2022) mace-off_medium Kovács et al. (2023)
mace Batatia et al. (2022) mace-off_small Kovács et al. (2023)
lmlp Eckhoff and Reiher (2023) <user-defined> Eckhoff and Reiher (2023)
II.7 AutoCAS
II.7.1 Field of application
Highly accurate electronic structure calculations require one to take electron correlation into consideration. Even though closed-shell molecules can be routinely and reliable calculated black-box coupled-cluster modelsBartlett (2024), strongly correlated systems require an active orbital space for a multi-configurational description with subsequent treatment of the dynamic correlation. Since the manual selection of an active space requires chemical intuition and experience, this step is often time consuming, difficult to reproduce, and hardly possible for non-prototypical complex system such as the one in Ref. 132. Especially for methods which can handle large active spaces of more than 100 orbitals, like DMRGWhite (1992, 1993); Baiardi and Reiher (2020) and full configuration interaction quantum Monte CarloBooth, Thom, and Alavi (2009); Cleland, Booth, and Alavi (2010), a manual selection is neither feasible nor reliable. Although many approachesPulay and Hamilton (1988); Bofill and Pulay (1989); Tishchenko, Zheng, and Truhlar (2008); Bao et al. (2016, 2017); Sayfutyarova et al. (2017); Bao et al. (2018); Khedkar and Roemelt (2019); Sayfutyarova and Hammes-Schiffer (2019); Jeong et al. (2020); Li, Gagliardi, and Truhlar (2020); Gieseking (2021); Golub et al. (2021); Khedkar and Roemelt (2021); Lei, Suo, and Liu (2021); Levine et al. (2021); Oakley, Gagliardi, and Truhlar (2021); King and Gagliardi (2021); Weser et al. (2022); Casanova (2022); Cheng, Xie, and Ma (2022); King et al. (2022); Kaufold et al. (2023); Golub et al. (2023), have been devised to address this problem (at least in parts), we have demonstrate in 2016 a viable, reliable, and truly fully automated selection approach based on quantum information measures, the autoCAS approach Stein, von Burg, and Reiher (2016); Stein and Reiher (2016); Stein2017; Stein and Reiher (2019); Bensberg and Reiher (2023b), a development, which has not been paralleled so far.
The autoCAS algorithmStein and Reiher (2016, 2019) exploits the fact that DMRG can converge a qualitatively correct wave function with less effort than what would be required for a converged energy.Moritz and Reiher (2007) Electron correlation encoded in this (approximate) wave function can be efficiently dissected in terms of orbital entanglement measuresBoguslawski et al. (2012). Hence an approximate DMRG calculation for all valence orbitals (including also double shells such as $4d$ orbitals for $3d$ transition metal atoms) delivers approximate, but qualitatively correct one- and two-orbital reduced density matrices, from which the single orbital entropies and the mutual informationLegeza and Sólyom (2003); Rissler, Noack, and White (2006); Legeza and Sólyom (2006) are evaluated. Based the single-orbital entropies, autoCAS defines the orbitals for an active space in a following fully converged CASSCF or DMRG calculation, optionally with subsequent dynamic correlation treatment. Even though the autoCAS algorithm is computationally more expensive than other active space selection algorithms, it is based on a multi-configurational description of the wave function and it is therefore not biased toward a single configuration, by contrast to other approaches. Moreover, since the approximate initial DMRG calculation relies on a rather small bond dimension and only a few sweeps to be efficient, the final fully converged calculation allows one to assess whether the initial settings were reliable enough for the active space selection. Should a discrepancy be detected because of truly strong correlations in the electronic structure, then one can easily restart the whole procedure with an increased bond dimension.
II.7.2 I/O
The autoCAS program packageBensberg et al. (2023f) provides different workflows for black-box active space calculations. Every workflow starts with a Hartree–Fock calculation, a subsequent active space selection, and a final calculation with the automatically selected active space. For small or test calculations the command-line interface provides the options to specify a basis set and a XYZ file containing the molecular structure. However, since the command-line interface is limited, production runs should use a YAML configuration file. Especially crucial options, such as the large active space protocolStein and Reiher (2019) and the required number of orbitals in each sub-space, can be specified there.
A user-friendly option to interact with autoCAS is provided by the graphical user interface HeronBensberg et al. (2022). Heron does not only provide more control over autoCAS, which is especially helpful for an unexperienced user, it also features a built-in orbital viewer. The true power of autoCAS, however, comes by including the active space selection in custom workflows through the API.
II.7.3 API
In the design of autoCAS, we focused on a strict separation of the algorithms and the electronic structure backend. An example autoCAS script could have the following form:
1from scine_autocas import Autocas
2from scine_autocas.autocas_utils.molecule import Molecule
3from scine_autocas.interfaces.molcas import Molcas
4
5xyz_file = "/path/to/molecule.xyz"
6# Create a molecule
7mol = Molecule(xyz_file)
8ac = Autocas(mol)
9# Initialize interface
10molcas = Molcas([mol])
11molcas.settings.basis_set = "cc-pVDZ"
12molcas.settings.xyz_file = xyz_file
13
14# Basic Workflow
15occ, ind = ac.make_initial_active_space()
16# HF calculation
17molcas.calculate()
18# CAS calculation
19e, s1, s2, mi = molcas.calculate(occ, ind)
20occ, ind = ac.get_active_space(occ, s1)
21# Final calculation
22e, s1, s2, mi = molcas.calculate(occ, ind)
Listing 6: Minimal autoCAS workflow with the OpenMolcas interface.
The Molecule class stores all molecule-specific values, such as spin and charge. It can be initialized through a XYZ file or a list of atoms (strings with element labels). The Autocas class is responsible for the handling of active spaces, either the valence space (number of occupied and virtual orbitals around the Fermi vacuum) or the selected active space based on single-orbital entropies. These functions return 2 lists. One list with the orbital indices and one list storing an occupation number for each orbital with a possible value of 2, 1 or 0, representing a doubly occupied, singly occupied or virtual orbital, respectively. Hence, the total number of electrons in an active space can be determined by summation of the entries in the occupation number list. Even though every interface in autoCAS inherits from a base class, customized workflows only need to provide a list with single-orbital entropies in the same order as the index and occupation lists.
II.8 ReaDuct
II.8.1 Field of application
Chemical reaction exploration relies on the determination of minima and first-order saddle points on potential energy surfaces. Minima correspond to stable molecular structures, while first-order saddle points represent transition state (TS) structures, which connect minima in reaction valleys. The tasks required to locate and connect these points on a PES can be performed by the SCINE module ReaDuctVaucher and Reiher (2018); Bensberg et al. (2023e). ReaDuct is designed in such a way that the execution of a task is entirely independent of the method characterizing the PES and providing the necessary quantum chemical properties, such as energies and nuclear gradients.
A typical workflow for discovering new reactions involves, starting from one (single-ended) or two (double-ended) minima, an initial transition state search task followed by a transition state optimization task. The optimized TS structure can then be associated to an elementary step by an intrinsic reaction coordinate (IRC) task, whose resulting endpoints can be optimized with a structure optimization.
Each task offers various options and is highly adaptable to the chemical problem at hand. A comprehensive and detailed description of all options can be found in the manual of the module. A few key options are listed in the following. For instance, given one minimum structure, TS guess structures to another minimum can be found with a transition state search task employing our Newton Trajectory Algorithm 1Unsleber, Grimmel, and Reiher (2022), Algorithm 2Unsleber, Grimmel, and Reiher (2022), and AFIR algorithm developed by the Maeda groupMaeda and Morokuma (2010, 2011); Maeda, Taketsugu, and Morokuma (2014). For the TS structure optimization, Hessian-based algorithms such as Eigenvector followingBanerjee et al. (1985); Schlegel (2011), an approach proposed by BofillBofill (1994) or a solely gradient-based algorithmHenkelman and Jónsson (1999); Kästner and Sherwood (2008); Shang and Liu (2010) are available. To connect a TS structure to reaction valleys, an IRC calculation along the eigenvector of the lowest eigenvalue of the Hessian, employing a steepest decent optimizer, can be executed. The resulting endpoints can be further optimized with various optimizers, which either explicitly calculate or approximate the Hessian, to obtain the local minima on the PES connected by the previously determined TS structure.
II.8.2 I/O
All functionality of ReaDuct can be accessed via an input file following the YAML syntax. The input file comprises two blocks: one defining multiple systems with their nuclear coordinates and the method defining the PES. The second block comprises the requested tasks with the relevant settings. Multiple tasks can be added and are executed after each other. The ReaDuct program is then simply run from the command line with
1readuct -c input.yaml
II.8.3 API
The ReaDuct module seamlessly integrates into workflows through its Python bindings. Executing tasks differs slightly from directly calling the ReaDuct binary. Instead of a single block for the system in an input file, one utilizes the Python bindings of our SCINE Utilities to create the corresponding Python object for the system (calculator).
Various tasks can then be executed through simple function calls, with the system (calculator) as an argument. The function returns a boolean (indicating whether the task was successful or not) and a Python dictionary containing the calculator with its results:
1import scine_utilities as utils
2import scine_readuct as readuct
3import scine_sparrow
4
5water_calculator = utils.core.load_system("h2o.xyz", "PM6")
6
7systems = {}
8systems["water"] = water_calculator
9systems, success = readuct.run_bond_order_task(systems, ["water"])
10results = systems["water"].get_results()
11print("Energy", results.energy)
Listing 7: Performing a bond order calculation task with ReaDuct.
The results Python object may have different attributes, such as the energy or the bond orders. Task-specific settings are provided as keyword arguments for the ReaDuct functions.
1optimization_settings = {"output": ["water_opt"], "optimizer": "BFGS"}
2systems, success = readuct.run_optimization_task(systems, ["water"], **optimization_settings)
3opt_results = systems["water_opt"].get_results()
4print("Opt. Energy", results.energy)
Listing 8: Performing a structure optimization task with ReaDuct.
This allows for sequential execution of various tasks, with results analyzed in a straightforward Python workflow. A prime example of such workflows is demonstrated in the numerous jobs within our SCINE Puffin module (see next section).
II.9 Puffin
II.9.1 Field of application
Puffin is a program that is launched in high numbers on high-performance computing infrastructure in order carry out composite workflows. The program is currently employed in, but not limited to, our exploration framework as it can encode any algorithm that requires no more than a set of atomistic input structures and an electronic structure model. Therefore, possible fields of application are high-throughput campaigns that carry out individual energy evaluations, optimization algorithms, or more complex workflows based on chained optimization algorithms such as reaction trials that combine reactive complex formation, Newton trajectory, single-ended transition state search, intrinsic reaction coordinate calculations, and additional checks to assess the bonding patterns, molecular charge and spin multiplicity of any found product(s).
To carry out these workflows, Puffin depends on other SCINE modules such as ReaDuct and possibly external quantum chemistry software. To simplify the setup of this large software stack, we built a so-called bootstrapping procedure through which a Puffin instance automatically downloads and compiles all required software. Large-scale deployments on high-throughput computing infrastructure are further simplified by running Puffin in a virtualized environment. For this, we currently support Docker and Apptainer images.
II.9.2 API
A new workflow can be defined by defining a new Job. A job is generally defined by its unique name that signals to a Puffin instance which workflow it must execute, and a run method that is carried out during execution in between general jobs, resource, and data management.
On top of that, Puffin automatically manages the different available resources, programs, and program versions so that it only carries out workflows appropriate to a specific instance and logs the employed programs. One computational resource that is difficult to estimate, especially in high-throughput settings, is the memory usage. Therefore, each instance receives a maximum memory requirement. During the workflow execution, Puffin monitors itself to stay below this limit and aborts the operation otherwise.
II.10 Database
II.10.1 Field of application
The database plays a central role in our automated explorations as it is not only the permanent data storage, but also the mean of communication between the software driving the exploration forward, SCINE ChemotonUnsleber, Grimmel, and Reiher (2022); Bensberg et al. (2023a), and the software executing individual calculations, SCINE PuffinBensberg et al. (2023c).
The sheer size of chemical reaction networks dictates to store the exploration progress in a database that can be scaled to millions of entries. Furthermore, the quantum chemical exploration of CRNs generates additional information about the electronic structure of individual chemical structures and minimum energy paths, which can be stored for subsequent campaigns to build data-driven modelsSimm and Reiher (2018); Gugler and Reiher (2022); Eckhoff and Reiher (2023).
In case the exploration is carried out autonomously, i.e., calculations are setup iteratively based on the already explored reaction space, it is desirable to employ the same database for both, carrying out calculations and storing the CRN. However, this requires the database to be able to handle thousands of simultaneous queries from the algorithms driving the exploration and the processes that carry out individual electronic structure calculations. Hence, we have developed a database schema for a MongoDBMongoDB Inc. database and a custom wrapper to interact with the databaseBensberg et al. (2023g). The wrapper ensures that every interaction with the database is synchronized with its current state. This guarantees that multiple processes can act on the current database simultaneously.
MongoDB databases consist of separate documents. In our database schema, different types of documents are defined, which are stored in collections of identical document types. The database encodes the CRN by linking several documents across multiple collections. Additionally, information about the carried out calculations and their results are stored in separate collections. The different document types are defined as:
  • Structure: a chemical structure defined by Cartesian coordinates of atoms, molecular charge, and spin multiplicity, representing a point on a Born–Oppenheimer potential energy surface
  • Property: a quantum chemical property of a specific structure
  • Elementary Step: a transformation of a set of structures to a different set of structures with an allowed intersection of the two sets
  • Calculation: a sequence of individual quantum chemical calculations on a set of structures that can generate new structures, properties, and elementary steps
  • Compound: a collection of structures that have an identical element composition, molecular charge, spin multiplicity, connectivity, and local arrangement around each nucleus. The latter is determined by the shape fitting and ranking algorithms in SCINE MolassemblerSobez and Reiher (2020); Bensberg et al. (2023b)
  • Flask: identical to Compound, but the contained structures are not required to be connected in one single graph
  • Reaction: a collection of elementary steps that transform structures belonging to identical compounds or flasks
We refer to Ref. 178 for more details on the individual definitions. The relations between all documents in the SCINE Database are illustrated in Fig. 4.
Figure 4: Schema of the SCINE Database in a MongoDB format. The document symbols show the different types of documents, which are stored in one collection each. The relationships between the documents are facilitated through cross references by unique identifiers in each document.
II.10.2 API
The SCINE Database is built around the general MongoDB database. In order to run a SCINE Database a MongoDB instance must be started and be addressable with an IP address, port, database name, and optional authentication credentials. Since SCINE Database is a C++ library, the database can also be addressed in C++ programs. However, our library also features Python bindings with which we showcase its capabilities here. We connect to the database with the Manager class:
1import scine_database as db
2
3manager = db.Manager()
4manager.set_credentials(db.Credentials("localhost", 27017, "database_name"))
5# The following line will fail if you don’t have
6# a database running or incorrect credentials
7manager.connect()
8# If you are creating a database from scratch
9# you can run
10#    manager.init()
11# to create the collections accessed in the
12# following
13calculations = manager.get_collection("calculations")
14compounds = manager.get_collection("compounds")
15flasks = manager.get_collection("flasks")
16elementary_steps = manager.get_collection("elementary_steps")
17reactions = manager.get_collection("reactions")
18properties = manager.get_collection("properties")
19structures = manager.get_collection("structures")
Listing 9: Connecting to the SCINE database and accessing different collections.
In order to retrieve information from the database, the collections objects hold the querying methods query_x, iterate_x, and count, x being a placeholder for the applied collection such as query_calculations or query_structures. The querying methods (cf. Listing 10) can take any valid MongoDB query, which are given as documents that specify which keys of the database documents should fulfill certain criteria, e.g., equality of values. More complex matches are achieved with operators, which start with \$, and multiple queries can be combined with logical operators such as “and”/“or”.
1from json import dumps
2
3# Count everything in the calculations
4# collection
5print("Total nr. of calculations:", calculations.count("{}"))
6# Count failed calculations
7print("Nr. of failed calculations:", calculations.count(dumps({"status": "failed"})))
8# Get calculation IDs of all successful DFT
9# calculations
10selection = {"\$and": [
11    {"status": {"\$in": ["complete", "analyzed"]}},
12    {"model.method_family": "DFT"}
13  ]
14}
Listing 10: Overview of the basic querying method with the SCINE database.
The underlying electronic structure methods can generally be summarized in the Model objects, which can be directly queried for. The results of the queries can be built directly into loops to aggregate information. For example, one can collect the Cartesian coordinates of all structures that were optimized with GFN2-xTB, but have already received a single-point energy obtained with an exchange-correlation density functional:
1from scine_database.queries import model_query, optimized_labels
2
3structure_model = db.Model("GFN2", "GFN2", "")
4structure_model.solvent = "water"
5structure_model.solvation = "any"
6energy_model = db.Model("DFT", "PBE-D3BJ", "def2-SVP")
7
8results = {}
9for structure in structures.query_structures(dumps(
10    {"\$and": [
11        {"exploration_disabled": False},
12        {"label": {"\$in": optimized_labels()}}
13    ] + model_query(structure_model)
14    }
15)):
16    prop_ids = structure.query_properties({
17        "electronic_energy",
18        energy_model,
19        properties
20    })
21    if prop_ids:
22        atoms = structure.get_atoms()
23        # Take most recent energy
24        energy_prop = db.NumberProperty(prop_ids[-1], properties)
25        results[str(structure.id())] = {
26            "elements": atoms.elements,
27            "coords": atoms.positions,
28            "energy": energy_prop.get_data()
29        }
Listing 11: Example for querying all structures of a certain model and aggregating their coordinates based on additional criteria and their electronic energy with a different model.
More detailed queries, such as the spread of the forward barriers of all elementary steps within a reaction, might require aggregated information that is not directly stored in the database, but can be calculated on the fly. In addition to the Python bindings, the SCINE Database library also features Python functions that carry out such aggregated queries which are required in reaction network explorations similar to the one in Listing 11.
II.11 Chemoton
II.11.1 Field of Application
The exhaustive elucidation of reaction mechanisms demands mapping all competitive elementary steps into a chemical reaction network, for which various automated approaches have been devisedSameera, Maeda, and Morokuma (2016); Dewyer and Zimmerman (2017); Simm, Vaucher, and Reiher (2019); Unsleber and Reiher (2020); Steiner and Reiher (2022); Baiardi et al. (2022); Ismail, Majerus, and Habershon (2022); Wen et al. (2023); Margraf et al. (2023). Our SCINE Chemoton module facilitates the exploration of CRNs in a fully automated fashion based on first principlesBergeler et al. (2015); Simm and Reiher (2018); Unsleber, Grimmel, and Reiher (2022). When provided with at least one starting molecule, Chemoton systematically discovers all possible reactions originating from this molecule and its successors. This process thereby reveals the entire reaction network. The modularity of Chemoton enables this comprehensive explorationUnsleber, Grimmel, and Reiher (2022).
An automated exploration requires several distinct tasks, each carried out independently and in parallel. The execution must be identical for all tasks, with tasks requiring the same architecture but addressing different facets of the automated exploration. Within our framework, we designate the execution component as the Engine and the task-specific aspect as the Gear. Each Engine is paired with a Gear and can be run individually.
Figure 5: Overview of Chemoton gears required for a chemical reaction network exploration.
Key gears of the Chemoton workflow are depicted in Fig. 5 and will be briefly explained in the following. The top-right gear in Fig. 5, the Kinetics gear, determines the molecules (aggregates) allowed for a reaction search. Moving clockwise, the Elementary Step Search gear probes one or two chemical species for a reaction, possibly in various orientations and along various reaction coordinates. The Aggregate Validation gear ensures that newly found molecules are local minima on their PES, sorting them into collections called “aggregates” of molecules based on multiplicity, charge, and bonding pattern. The latter is analyzed with Molassembler (see Section II.1 for details). Similarly, the Reaction Sorting gear collects reactions connecting the same aggregates (see Section II.10 for the nomenclature employed). To distribute the available resources based on each gear’s needs, the Job Scheduling gear determines the order in which calculations should be executed. This guarantees a seamless exploration and full usage of resources. The Refinement gear allows for recalculating certain properties, such as the energy of a molecule, e.g., with a different electronic structure method.
The software structure presented so far is strongly tied to autonomous, indefinitely running, independently working processes, which is well suited for exhaustive explorations based on defined parameters for the complete exploration. However, if the chemical system or reaction mechanism is highly complex and does not lend itself easily to exhaustive explorations, a linear style of exploration can establish important nodes in the reaction network before expanding then on the paths obtained with automated approaches. Such a linear exploration is enabled by the top-level Steering Wheel frameworkSteiner and Reiher (2023) that combines several gears into discrete exploration steps. The exploration steps are either a Network Expansion that adds information to the reaction network or a Selection Step that defines a subset of the network. These must be applied in an alternating fashion which allows one to systematically steer an exploration into certain directions of chemical reaction space.
All calculations created by the running gears and required for an exploration are written to a database (Section II.10). Once they are in the database, they are simply jobs for our calculation backend Puffin (Section II.9), responsible for running the jobs and writing their results into the database. For setting up new calculations and expanding the exploration, the running Chemoton gears are constantly checking on these results. Hence, all of the results and output created by running a chemical network exploration with Chemoton is stored in a database and chemical insights can be gained by querying this database.
The resulting networks are often too extensive and interconnected to be easily interpreted by a human being. Therefore, we developed Pathfinder within Chemoton to facilitate navigation and analysis of such networksTürtscher and Reiher (2023). Pathfinder converts the network stored in the database (Section II.10) to a graph, encoding both stoichiometric and kinetic information in the edges. Given user-defined starting conditions, the aggregates in the graph have varying likelihood of being encountered based on their accessibility from a starting aggregate. The edges of the graph are then updated with this information, ensuring that unaccessible aggregates are disfavored in a simple shortest path search. In such a search, source and target vertices can be any vertices in the graph, providing a route in terms of reaction sequences for how a certain product is formed from a given starting aggregate under the specified starting conditions. Altering the starting conditions can lead to different shortest routes for a certain source–target vertex pair, allowing the investigation of various conditions. Overall, Pathfinder facilitates efficient navigation through and analysis of complex chemical reaction networks efficiently, providing intuitive insights for chemists in the form of reaction sequences and profiles.
II.11.2 API
Due to the scale of automated explorations, their setup depends strongly on the specific use case and desired properties. The most-approachable entry point is within our graphical user interface, which allows one to construct individual gears and offers pop-up menus for all their options. More experienced users can setup the individual gears in a small Python script as shown in Listing 12. The Chemoton main script serves as template for such scripts.
1from scine_chemoton.gears.thermo import BasicThermoDataCompletion
2from scine_chemoton.gears.scheduler import Scheduler
3from scine_chemoton.engine import Engine
4from scine_database import Credentials
5
6credentials = Credentials("localhost", 27017, "database_name")
7# Hessian calculations
8thermo_gear = BasicThermoDataCompletion()
9thermo_engine = Engine(credentials, fork=False)
10thermo_engine.set_gear(thermo_gear)
11thermo_engine.run(single=True)
12# Handle calculation scheduling
13schedule_gear = Scheduler()
14schedule_gear.options.job_counts["scine_hessian"] = 100_000
15schedule_engine = Engine(credentials, fork=False)
16schedule_engine.set_gear(schedule_gear)
17schedule_engine.run(single=True)
Listing 12: Example for running a Chemoton engine with a specific gear.
A steered exploration can be carried out in a Python environment by constructing the linear exploration protocol. The example in Listing 13 sets up an exploration that adds two compounds to the database, creates conformers, selects the lowest energy conformers in each cluster based on k-means clustering, and samples possible reactions between the compounds. Later, the protocol is enhanced with a selection of all found products with updated reactive site rules and then only unimolecular rearrangements are sampled. Due to the linear execution logic of the Steering Wheel, it does not matter whether the two last steps were added later or from the beginning of the exploration run as both procedures produce identical results.
1from time import sleep
2from scine_chemoton.steering_wheel import SteeringWheel
3from scine_chemoton.steering_wheel.selections import AllCompoundsSelection
4from scine_chemoton.steering_wheel.selections.input_selections import FileInputSelection
5from scine_chemoton.steering_wheel.selections.conformers import LowestEnergyConformerPerClusterSelection
6from scine_chemoton.steering_wheel.network_expansions.basics import SimpleOptimization
7from scine_chemoton.steering_wheel.network_expansions.conformers import ConformerCreation
8from scine_chemoton.steering_wheel.network_expansions.reactions import Association
9from scine_chemoton.gears.elementary_steps.reaction_rules.reaction_rule_library import (
10    DefaultOrganicChemistry,
11    SimpleDistanceRule,
12)
13from scine_chemoton.gears.elementary_steps.reactive_site_filters import AtomRuleBasedFilter
14import scine_database as db
15
16model = db.Model("GFN2", "GFN2", "")
17protocol = [
18    FileInputSelection(
19        model,
20        [
21        # Structure file, charge, spin mult.
22        ["molecule_A.xyz", 0, 1],
23        ["molecule_B.xyz", 0, 1]
24    ]),
25    SimpleOptimization(model, status_cycle_time=1),
26    AllCompoundsSelection(model),
27    ConformerCreation(model),
28    LowestEnergyConformerPerClusterSelection(
29        model,
30        n_clusters=5,  # k-means
31        additional_reactive_site_filters=AtomRuleBasedFilter(
32            DefaultOrganicChemistry()
33        )
34    ),
35    Association(
36        model,
37        max_bond_associations=2,
38        max_bond_dissociations=1
39    )
40]
41wheel = SteeringWheel(
42    db.Credentials("localhost", 27017, "database_name"),
43    protocol
44)
45wheel.run()  # runs non-blocking
46sleep(1)
47# One can query progress along the way:
48print(wheel.get_status_report())
49print("will run until all finished or stopped")
50print("will only finish if calculations are executed")
51try:
52    while wheel.is_running():
53        sleep(1)
54except KeyboardInterrupt:
55    wheel.terminate()
56wheel.save()
Listing 13: Carrying out a steered reaction exploration with Chemoton.
Due to the top-level approach and flexibility of the Steering Wheel, the decision on the exploration protocol requiresva priori either a hypothesis about the possible reaction mechanism or a good overview of the status of the exploration in order to adapt the exploration strategy on the fly. Because the first requirement is rare, we have built the framework around the second requirement by integrating it into our graphical user interface Heron. This allows one to access the current exploration status and preview the effect of potential further exploration steps.
II.12 KiNetX
II.12.1 Field of application
Microkinetic modeling simulations are crucial to connect the microscopic description of a chemical reaction as a detailed reaction network to macroscopic observations such as yields by predicting the concentration fluxes through the network. Microkinetic modeling simulations can be run through the SCINE module KiNetXBensberg et al. (2023d). Moreover, such simulations can automatically steer the reaction network explorationBensberg and Reiher (2023a) by interweaving them directly in a rolling exploration. KiNetX is interfaced with the library SundialsHindmarsh et al. (2005); Gardner et al. (2022), which provides efficient integration routines for the differential equations in the microkinetic modeling. As an alternative to KiNetX, SCINE provides an interface to the program Reaction Mechanism Simulator (RMS)Johnson, McGill, and Green (2022); Johnson et al. (2023) for microkinetic modeling. To predict which reactions and compounds in a reaction network dominate the overall kinetics and therefore must be described accurately, the SCINE framework provides sensitivity analysis approaches for the RMS microkinetic modeling through the library SALibHerman and Usher (2017); Iwanaga, Usher, and Herman (2022). The sensitivity analysis can also be interwoven with a rolling exploration to identify the key reactions and compounds and refine their descriptions by reoptimizing structures or recalculating electronic energies with more accurate electronic structure modelsBensberg and Reiher (2023c).
II.12.2 I/O
KiNetX assumes constant volume and temperature for the microkinetic modeling. It requires only rate constants, reactive species, reactions, start concentrations, and simulation time or number of time steps as input. During the simulation, it prints the concentrations for batches of time steps to the screen. The calculation input must be provided through its Python API.
II.12.3 API
KiNetX can be easily integrated into automated workflows through its Python bindings. First, the forward/backward rate constants (k_fs/k_bs), reaction stoichiometry (s_fs/s_bs), and the compounds must be encoded in a Python object that represents the reaction network. This is accomplished by successively adding compounds and reactions to a NetworkBuilder object:
1import scine_kinetx as kx
2network_builder = kx.NetworkBuilder()
3for mass, name in zip(compound_masses,
4                      compound_names):
5    network_builder.add_compound(
6        mass, name)
7for k_f, k_b, s_f, s_b in zip(k_fs, k_bs,
8                              s_fs, s_bs):
9    network_builder.add_reaction(k_f, k_b,
10                                 s_f, s_b)
Listing 14: Encoding the reaction network for KiNetX (for the sake of brevity, this code listing cannot be executed as such, since variables such as compound_masses are not defined).
Then, the microkinetic model can be integrated by calling the integrate function on the network built by the NetworkBuilder with the start concentrations c_start, the start time t_start, initial time step dt, final time t_max, and integrator choice:
1network = network_builder.generate()
2result = kx.integrate(
3    network, c_start,
4    t_start, dt,
5    kx.Integrator.cash_karp_5,
6    integrateByTime=True,
7    maxTime=t_max)
Listing 15: Solving the microkinetic model with KiNetX.
As a result, the integrate function returns the final and maximum concentrations, the integrated absolute concentration fluxesBensberg and Reiher (2023a) through the compounds and reactions.
III Conclusions
Here, we presented the modules of the SCINE project that provide a framework for quantum chemical calculations on isolated and embedded molecular structures that are to be arranged in some conceptual framework. Typical frameworks that provide contextual dependence for these structures are chemical reaction networks or high-throughput virtual screening campaigns in an attempt to optimize or design certain properties (such as linear or nonlinear optical properties or specific reactivity patterns). Although the focus of our work has been mostly on the former application area so far, the latter is easily realizable within the SCINE framework. Modules that produce data on molecular structures can be easily combined to accomplish high-level tasks (as highlighted by the steering through a reaction mechanism during reaction network exploration with the Steering WheelSteiner and Reiher (2023)). Data is stored in and recovered from a central database structure. The Molassembler module produces unique graph information for the identification of all kinds of molecular structure, which is key for the identification, storage, and exploitation of Cartesian coordinates as molecules. The high degree of modularity and interoperability of the SCINE programs has facilitated the development of the cloud-based reaction mechanism screening workflow AutoRXNUnsleber et al. (2023). In general, the SCINE project has established a platform of high usability for further developments in the field of high-throughput computational campaigns.
Acknowledgements.
This publication was created as part of NCCR Catalysis, a National Centre of Competence in Research funded by the Swiss National Science Foundation (grant number 180544). M. E. gratefully acknowledges an ETH Zurich Postdoctoral Fellowship. M. S. gratefully acknowledges a Swiss Government Excellence Scholarship for Foreign Scholars and Artists. M. M. gratefully acknowledges financial support through ETH Research Grant ETH-43 20-2.
Data Availability Statement
Data sharing is not applicable to this article as no new data were created or analyzed in this study.